Find the magnitude of an electron's electric field if the field itself is 50cm away from it.

1
Expert's answer
2022-03-06T15:16:13-0500

The electric field

"E=k\frac{q}{r^2}\\
=9*10^9*\frac{1.6*10^{-19}}{0.5^2}=5.76*10^{-9}\:\rm V/m"